# High-entropy Advantage in Neural Networks’ Generalizability

### v1

@@ Abstract
While the 2024 Nobel Prize in Physics ignites a worldwide discussion on the origins of neural networks and their foundational links to physics, modern machine learning research predominantly focuses on computational and algorithmic advancements, overlooking a picture of physics. Here we introduce the concept of entropy into neural networks by reconceptualizing them as hypothetical physical systems where each parameter is a non-interacting ’particle’ within a one-dimensional space. By employing a Wang-Landau algorithms, we construct the neural networks’ (with up to 1 million parameters) entropy landscapes as functions of training loss and test accuracy (or loss) across four distinct machine learning tasks, including arithmetic question, real-world tabular data, image recognition, and language modeling. Our results reveal the existence of entropy advantage, where the high-entropy states generally outperform the states reached via classical training optimizer like stochastic gradient descent. We also find this advantage is more pronounced in narrower networks, indicating a need of different training optimizers tailored to different sizes of neural networks.
@@
Neural Networks (NNs) have shown remarkable success in solving a wide range of machine learning problems, from image recognition to natural-language processing. [1, 2, 3, 4] Despite their success, applications can face the problem of overfitting [5, 6, 7, 8], which occurs when an NN learns not only the underlying patterns in the training data, but also the noise and random fluctuations. Consequently, the NN could perform exceptionally well on the training data but fail to generalize to new, unseen data. The ability to generalize is a critical metric to evaluate their usefulness in the real-world applications, making generalizability one of the central topics in the field. [9, 6, 8] Modern NNs are usually overparameterized, i.e., the number of parameters exceeds the number of training data points.[10, 11] Intuitively, an overparamterized model can easily overfit the training data, akin to how a high-order polynomial fit generally fails to capture the underlying trend of the data. In fact, today’s NNs have sufficient capacity to memorize the entire training set, as evidenced by their ability to easily fit data with random labels.[6] However, despite their massive size, NNs have achieved state-of-art generalization performance in many different machine-learning tasks.[12, 13, 14] Such surprising generalization has sparked extensive research to understand its underlying mechanism.
@@
Analogous to the representation of all degrees of freedom in a physical system as a point in high-dimensional phase space, all trainable parameters in a neural network can likewise be depicted as a point in a similarly high-dimensional parameter space. For an over-parameterized neural network, there are infinitely many points in the parameter space that can perfectly fit the training data. Some of the points generalize well, while others do not. Therefore, many studies on generalizability looked for its correlation with certain characteristics in the parameter space, including sharpness, the VC dimension, Rademacher complexity, weight norms, and intrinsic dimension[15, 16, 17, 18]. Unfortunately, this line of research has not yet generated convincing conclusions,[16] as the best-performed measure, sharpness, has been challenged to be sufficient enough to determine the generalization by itself.[19, 20] There are also studies attributing the generalizability to different regularization methods,[5, 21, 22, 6] which are protocols that bias the training trajectory toward some specific parts of the parameter space that have been empirically found to generalize well. Although empirical studies suggest that regularization can help generalization, many modern NNs can still generalize well without these methods, excluding them as the fundamental mechanism of generalization.[6]
@@
Here we investigate the mystery of NN generalizability from a physics perspective and show that among all states that fits the training data well, the highest-entropy ones are highly generalizable. Since entropy is the logarithm of the parameter-space volume, our results demonstrate that generalizable well-trained states occupy a larger part of the parameter space than its non-generalizable analog. This explains why training algorithms can find generalizable states, even when we do not actively apply regularization methods. In fact, our results show that max entropy states often generalize even better than states trained with stochastic gradient descent (SGD). In other words, SGD usually biases toward slightly less generalizable parameters, but the generalizable part of the parameter space is so massive that SGD-trained states can generalize well anyway. We intentionally choose to compare with SGD training rather than other popular optimizers like Adam, because it has been shown previously that SGD can usually realize better generalization in neural networks training. [23, 24, 25] We examined the entropy-generalizability relations in four distinct machine learning tasks, including arithmetic question, real-world tabular data, image recognition, and language modeling. Our results demonstrate that in all four tasks, the generalizability of high-entropy states are better or at least comparable with the states reached via the classical training optimizer like stochastic gradient descent (SGD). We refer to this as the high-entropy advantage in this paper. To better understand the condition of this advantage, we further studied its correlation with the network width. Our results indicate that wider NNs tend to have a smaller high-entropy advantage for the same task (e.g. reaching the limit of generalization), which agrees with derivations under the infinite network width limit and the reported observations. [26]
@@ Results
We begun our investigation from small fully-connected neural networks (FNNs) on a relatively simple arithmetic task: learning a predefined equation of binary classification. Here we constructed a synthetic ’spiral’ dataset, where every data point is a point on 2-D space with a color (brown or green) determined by its coordinates. More details including the equation used and figure of the generated dataset are presented in the Supplementary Materials. We generated 20 green points and 20 brown points randomly, forming two spirals. Our task is to predict whether a point is brown or green given its coordination. Therefore, this is a binary classification problem with 2 input features (horizontal and vertical coordinates).
We used Wang-Landau Monte Carlo (WLMC) method to sample the entropy landscape, $S(L_{train},A_{test})$, as a function of the log-scaled train loss, $\ln{(L_{train})}$, and test accuracy, $A_{test}$. The log-scaled train loss allows us to better sample the low $L_{train}$ states where the model converges after learning. In Fig. 1 (a), we presented the entropy landscape for a $4$ layers $\times$ $8$ neurons FNN (249 trainable parameters). Color gradient represents the entropy gradient, where warm color represents high entropy state. The WLMC method (and the Wang-Landau Molecualr Dynamics method we will employ later) calculates the entropy with an algorithm-dependent zero point; thus, the absolute value of the entropy is meaningless, but the entropy difference between any two points in Fig. 1 (a) is meaningful. For each given $\ln(L_{train})$, we can calculate the corresponding equilibrium test accuracy using Equation 1:
@@
We then performed the classical training via the SGD optimizer for 100 times and collected the $L_{train}$ versus $A_{test}$ trajectories. For each given $L_{train}$, we calculated the mean $A_{test}$, which is plotted as the black curve in Figure 1 (a). For simplicity, we refer it as the SGD accuracy in this paper. We can see that when $L_{train}$ reaches a low level ($\ln{L_{train}}\approx-0.5$), the equilibrium accuracy increases rapidly with the decrease of $L_{train}$ and gets saturated when $\ln{L_{train}}\approx-3$. In this regime, we observed a supremacy of the equilibrium accuracy over the SGD accuracy for each given $\ln{L_{train}}$, which indicates the existence of high-entropy advantage. Another interesting thing we noticed here is that when the training loss is at a high level ($\ln{L_{train}}>0$), the equilibrium accuracy is around 50%. This also agrees with our intuition, because when FNN has a high $L_{train}$, it generally suggests that the model does not learn anything yet but just making random guess, which should has an accuracy 50% for the binary classification.
We followed the same protocol and further tested on 3 additional FNN sizes and 2 different training time (see more details in Supplementary Materials). Results are presented in Figure 1 (b). Each point represents one training condition, and the black dashed line represents the level where equilibrium accuracy equals the SGD accuracy. Here we used the best SGD accuracy observed (i.e. the best-possible early stopping) and then looked for its corresponding equilibrium accuracy at the same train loss level. For all the 8 experiments conducted, the equilibrium accuracy outperforms the SGD accuracy with a large margin, further verifying that the high-entropy state can provide better generalization. The smallest FNN tested here ($3$ layers $\times$ $6$ neurons) has 109 trainable parameters, which is still much larger than the number of training data points, 40, guaranteeing the overparamterized nature of all models.
Figure 1: Entropy analysis via WLMC for arithmetic question: (a) Entropy landscape as a function of $ln(\mbox{train loss})$ and test accuracy. Color gradient represents the entropy gradient. (b) Equilbirum test accuracy versus the corresponding SGD test accuracy for $8$ different NNs.
@@
We started from the House Price dataset[27] on Kaggle website[28], where the task is to predict the the house price from its descriptors like number of bedrooms and house age. For the sake of computational cost, here we only used the original training dataset which consists of 1460 houses. Each house has 79 descriptors (excluding the house ID) and we performed one-hot encoding for all categorical descriptors, yielding a total of 331 descriptors / features for each data point eventually. We randomly select 50% data as the train set, with the remaining data reserved for test. A 2-layer FNN is used for this regression task, where the hidden layer has 20 neurons. Therefore, the final model has 6661 trainable parameters, making it overparamterized.
Following the similar protocol described above for the arithmetic question, we got the entropy landscape, $S(L_{train},L_{test})$, as a function of train loss, $L_{train}$, and test loss, $L_{test}$. Results are presented in Figure 2 (a), where both $L_{train}$ and $L_{test}$ are in log scale for better sampling the low loss regime. As in statistical physics, the probability distribution over different states (different locations in the plot) is sharply peaked around the entropy maximum, making the max-entropy state the thermodynamic equilibrium. We find out these max-entropy states at each $L_{train}$. We then performed the classical training via the SGD optimizer for 100 times and 3 different batch sizes, to make sure we have enough data points at each $L_{train}$ value. $L_{train}-L_{test}$ trajectories are collected and the mean $\ln{L_{test}}$ at each $\ln{L_{train}}$ are plotted as the black curve. The error bar for SGD training is smaller than the width of the black curve. Our results suggest that at a given $L_{train}$, the corresponding max entropy loss is significant lower than the test loss obtained via SGD training, suggesting the high-entropy advantage in NNs generalizability. More technical details, including data and hyperparameters tuning, are presented in Supplementary Materials.
Figure 2: Entropy landscape of (a) Kaggle Housing Price dataset; (b) MNIST Image dataset; (c) Polymer SMILES dataset. For the SGD training of all three models, the standard error at each train loss level is smaller than the width of the curve.
@@
The modern neural networks have demonstrated remarkable success in the realm of computer vision [29]. One of the fundamental datasets used in the computer vision benchmark is the MNIST dataset, which consists of images with handwritten digits. [30] Here we used it to further verify that the high-entropy advantage for the computer vision task. Condisdering many models have presented 100% accuracy on the test set of MNIST (see Leaderboard of Kaggle Digit Recognizer competition[31]), we strategically increase the difficulty of the MNIST task by drastically reducing the size of training dataset. This is done by randomly sampling 500 images from the original dataset and dividing them equally as the train and test sets, respectively. We refer this dataset as the selected MNIST dataset in this paper and more details can be found in the Supplementary materials. The original MNIST dataset has 50,000 images for training, which is 200 times larger than ours. Thus, we expect this constraint can significantly intensifies the task difficulties, so that we can better differentiate the potential advantage of the high-entropy states. We built a small convolutional neural network with 5 convolutional layers and one fully-connected layer. This network has 362 trainable parameters, which is larger than the training dataset (250 images) in the selected MNIST, making it an overparamterized model.
In Figure 2 (b), we presented the entropy landscape, $S(L_{train},A^{s}_{test})$, of this neural network on the selected MNIST dataset. Here $A^{s}_{test}$ is smoothed accuracy which is differentiable, as required by the WLMD algorithm (See Methods for details). We then find the max-entropy grid for each train loss on the landscape, which is again plotted as the magenta dots. Similar to the landscape in Figure 1 (a), we observed an increase of max entropy test accuracy when the train loss reaches a low level ($L_{train}\approx-1$) and it gets saturated when $\ln{L_{train}}\approx-4$. When the train loss is high, we observed a high-entropy band around the test accuracy of 0.1. This also meets our expectation because the digit recognition is a 10-class classification problem, so random guesses should have an accuracy of 10%. We then performed the classical training via the SGD optimizer for 200 times. (See more details in Supplementary Materials) The mean SGD accuracy are plotted as the black curve. Our results suggest that for a given train loss, the corresponding max entropy accuracy generally outperforms the SGD training, especially when the train loss is small ($\ln{L_{train}}<-2$).
In the Supplementary Materials, we further show the existence of the entropy advantage in a deeper neural network (10-layer ResNet [29]) on a selected CIFAR-10 dataset with $5,000$ images.
@@
Natural language modeling is another domain where the neural networks have consistently outperformed traditional machine learning methods. [2] Language modeling is also inherently differs from other machine learning tasks due to its reliance on semantic complexity and contextuality. [32] Therefore, we further extended our tests to a language modeling dataset. We utilized the TransPolymer model published recently, which achieves state-of-the-art performance in all ten different downstream tasks for polymer property prediction.[33]. TransPolymer is a BERT (Bidirectional Encoder Representations from Transformers) family model[3, 34] and is pretrained on rougly 5 millions polymer SMILES augmented from the Pl1M database[35]. This pretrained transformer-based language model is able to take the SMILES as input directly and generate a 768-dimension embedding for each given SMILES. The embedding is then fed into a regressor head, a fully-connected layer with SiLU activation function, and is regressed on different polymer properties. Here we choose the Egb Dataset[36, 33] to perform the entropy advantage experiment, which is the bandgap energy of bulk polymer and consists of 561 data points. This is the dataset where TransPolymer has the best performance (test $R^{2}=0.93$), therefore, we can verify whether entropy advantage still exists for such a well-learned task. 80% data is used for training and the remaining 20% data is reserved for testing, which is same as the way reported in the original TransPolymer paper[33].
For sampling efficiency, we follow the fintuning strategy of large foundation model[37, 38, 33], where the encoder embedding is fixed and only tuning the regressor head of the TransPolymer model. We also reduced the width of the regressor to 50 while keep the SiLU activation function. Therefore, the final model has $38501$ parameters, making it overparamterized. Entropy landscape of this language modeling task is presented in Figure 2 (c). As a regression problem, we found that the corresponding max-entropy loss at each $L_{train}$ is slightly lower or comparable to its SGD training analog, which suggests that high-entropy state can generalize well. In other words, even for a task that could be well-learned via the SGD training, there is still a high-entropy advantage.
@@
Now we have demonstrated high-entropy advantage in four distinct machine learning tasks, we will now study how is this advantage affected by the size of the neural network. It has been known that neural networks are equivalent to Gaussian processes in the limit of infinite width.[39, 40] Previous experiments also indicated that the performance of optimized neural networks converges to the corresponding Gaussian Process with increasing network width.[40] Consequently, it is reasonable to expect the high-entropy advantage could also change as a function of the network width.
To better evaluate this, we constructed a Spiral Regression Dataset (see Supplementary Materials for details) which consists of 500 data points and is divided equally for train and test sets. Using the same WLMD method, we constructed the entropy landscapes on NNs with four different widths, $W$, ranging from 30 to 1000. All four models are overparamerterized and the largest one has more than 1 million trainable parameters. Figure 3 presents the entropy landscapes of different networks and our results suggest that narrower networks possess more significant entropy advantage, which finally fades away when $W=1000$. Note here we used the Adam optimizer because the SGD optimizer performs much poorer. We also performed the same experiments on the Kaggle House price prediction, the selected MNIST image recognition, and the Polymer SMILES language modeling tasks and observed the similar trend: entropy advantage is more significant for narrower networks (See Figure 7 - 9 in the Supplementary Materials).

### v2

@@ Abstract
One of the central challenges in modern machine learning is understanding how neural networks generalize knowledge learned from training data to unseen test data. While numerous empirical techniques have been proposed to improve generalization, a theoretical understanding of the mechanism of generalization remains elusive. Here we introduce the concept of Boltzmann entropy into neural networks by re-conceptualizing such networks as hypothetical molecular systems where weights and biases are atomic coordinates, and the loss function is the potential energy. By employing molecular simulation algorithms, we compute entropy landscapes as functions of both training loss and test accuracy (or test loss), on networks with up to 1 million parameters, across four distinct machine learning tasks: arithmetic question, real-world tabular data, image recognition, and language modeling. Our results reveal the existence of high-entropy advantage, wherein high-entropy network states generally outperform those reached via conventional training techniques like stochastic gradient descent. This entropy advantage provides a thermodynamic explanation for neural network generalizability: the generalizable states occupy a larger part of the parameter space than its non-generalizable analog at low train loss. Furthermore, we find this advantage more pronounced in narrower neural networks, indicating a need for different training optimizers tailored to different sizes of networks.
@@ Introduction
Neural Networks (NNs) have shown remarkable success in solving a wide range of machine learning problems, from image recognition to natural-language processing. [1, 2, 3, 4] Despite their success, applications can face the problem of overfitting,[5, 6, 7, 8] which occurs when an NN learns not only the underlying patterns in the training data, but also the noise and random fluctuations. Consequently, the NN could perform exceptionally well on the training data but fail to generalize to new, unseen data. The ability to generalize is a critical metric to evaluate an NN’s usefulness in real-world applications, making generalizability one of the central topics in the field. [6, 8] Modern NNs are usually overparameterized, i.e., the number of parameters exceeds the number of training data points.[9, 10] Intuitively, an overparameterized model can easily overfit the training data, akin to how a high-order polynomial overfit a curve and fail to capture the underlying trend. In fact, today’s NNs have sufficient capacity to memorize the entire training set, as evidenced by their ability to easily fit data with random labels.[6] However, despite their massive size, NNs have achieved state-of-the-art generalization performance in many different machine-learning tasks.[11, 12, 13, 14] Such surprising generalizability has sparked extensive research to understand its underlying mechanism.
@@
Analogous to the representation of all degrees of freedom in a physical system as a point in high-dimensional phase space, each specific configuration of an NN, defined by the values of all its trainable parameters, can likewise be depicted as a point in a similarly high-dimensional parameter space. For an over-parameterized NN, there are many solutions in the parameter space that can fit the training data equally well.[15] Some of the solutions generalize well, while others do not. Therefore, many studies on generalizability looked for its correlation with certain characteristics in the parameter space, including sharpness, the VC dimension, Rademacher complexity, weight norms, and intrinsic dimension.[16, 17, 18, 19] Unfortunately, this line of research has not yet generated convincing conclusions,[17] as the best-performed measure, sharpness, has been challenged to be sufficient enough to determine the generalization by itself.[20, 15] There are also studies attributing the generalizability to different regularization methods,[5, 21, 6] which are protocols that bias the training trajectory toward some specific parts of the parameter space that have been empirically found to generalize well. Although empirical studies suggest that regularization can help generalization, many modern NNs can still generalize well without these methods, excluding them as the fundamental mechanism of generalization.[6]
@@
Here we investigate the mystery of NN generalizability from a physics perspective by conceptualizing it as a hypothetical molecular system. Each parameter in a NN (weights and biases) acts as an “atom” with its value representing its atomic coordinates. We then compute the Boltzmann entropy of this molecular system as a function of train loss and test performance (accuracy for classification or loss for regression). Our results show that among all states that fits the training data well, the highest-entropy ones are highly generalizable. Specifically, we examined the entropy-generalizability relations in four distinct machine learning tasks, including arithmetic question, real-world tabular data, image recognition, and language modeling. Our results demonstrate that in all four tasks, the generalizability of high-entropy states are better or at least comparable with the states reached via the classical training optimizer like stochastic gradient descent (SGD). We refer to this as the high-entropy advantage in this paper. We intentionally choose to compare with SGD training rather than other popular optimizers like Adam, because it has been shown empirically that SGD can usually realize better generalization in neural networks training. [22, 23, 24] Recent work also shows that the noise of SGD introduces an additional effective loss (implicit regularization) which favors flat solution in neural networks’ loss landscape and therefore can lead to better generalization.[25] To better understand the condition of high-entropy advantage, we further studied its correlation with the network width. Our results indicate that wider NNs tend to have a smaller high-entropy advantage for the same task (e.g. reaching the limit of generalization), which agrees with derivations and the reported observations under the infinite network width limit.[26]
@@ Results
We begun our investigation from small fully-connected neural networks (FNNs) on a relatively simple arithmetic task: learning a predefined equation of binary classification. Here we constructed a synthetic Spiral Classification Dataset, where every data entry is a point on 2-D space with a color (brown or green) determined by its coordinates. We generated 20 green points and 20 brown points randomly, forming two spirals. Our task is to predict a point’s color given its coordination. Therefore, this is a binary classification problem with 2 input features (horizontal and vertical coordinates). More details including the equation used and figure of this synthetic dataset are presented in the Supplementary Materials.
We used Wang-Landau Monte Carlo (WLMC) method to sample the entropy landscape, $S(L_{train},A_{test})$, as a function of the log-scaled train loss, $\ln{(L_{train})}$, and test accuracy, $A_{test}$. The log-scaled train loss allows us to better sample the low $L_{train}$ states, corresponding to trained NNs. In Fig. 1 (a), we present the entropy landscape for a ($3$ layers) $\times$ ($6$ neurons) FNN (116 trainable parameters). Warmer color represents higher entropy. The WLMC method (and the Wang-Landau Molecular Dynamics method we will employ later) calculates the entropy with an algorithm-dependent zero point; thus, the absolute value of the entropy lacks physical significance, but the entropy difference between any two points in Fig. 1 (a) carries physical meaning. For each given $\ln(L_{train})$, we can calculate the corresponding equilibrium test accuracy:
@@
We then performed classical training via the SGD optimizer for 100 times and collected the $L_{train}$ versus $A_{test}$ trajectories. For each given $L_{train}$, we calculated the mean $A_{test}$, which is plotted as the black curve in Figure 1 (a). For simplicity, we refer it as the SGD accuracy in this paper. We can see that when $L_{train}$ reaches a low level [$\ln(L_{train})\approx-0.5$)], the equilibrium accuracy increases rapidly with the decrease of $L_{train}$ and gets saturated when $\ln(L_{train})\approx-3$. In this regime, we observed a supremacy of the equilibrium accuracy over the SGD accuracy for each given $\ln(L_{train})$, which indicates the existence of high-entropy advantage. Another interesting observation is that when the training loss is at a high level ($\ln(L_{train})>0$), the equilibrium accuracy is around 50%. This also agrees with intuition, because when a FNN has a high $L_{train}$, the highest-entropy state of the model is just making random guesses, which results in 50% accuracy for binary classification.
We followed the same protocol and further tested on 3 additional FNN sizes and 2 different training time (total 8 experiments, see more details in Supplementary Materials). Results are presented in Figure 1 (b). For all experiments, the training loss at the end of SGD training corresponds to an equilibrium test accuracy that outperforms the SGD accuracy with a large margin, further verifying that the high-entropy state can provide better generalization. The smallest FNN tested here ($3$ layers $\times$ $6$ neurons) has 116 trainable parameters, which is still much larger than the number of training data points, 40, guaranteeing the overparamterized nature of all models.
Figure 1: Entropy analysis via WLMC for arithmetic question: (a) Entropy landscape as a function of $\ln(\mbox{train loss})$ and test accuracy. Color gradient represents the entropy gradient. (b) Equilbirum test accuracy versus the corresponding SGD test accuracy for $8$ different experiments.
@@
We started from the House Price dataset[27] on Kaggle website,[28] where the task is to predict the the house price from its descriptors like number of bedrooms and house age. For the sake of computational cost, here we only used the original training dataset which consists of 1460 houses. Each house has 79 descriptors (excluding the house ID) and we performed one-hot encoding for all categorical descriptors, yielding a total of 331 descriptors / features for each data point eventually. We randomly select 50% data as the training set, with the remaining data reserved for testing. A 2-layer FNN is used for this regression task, where the hidden layer has 20 neurons. Therefore, the final model has 6661 trainable parameters, making it overparamterized.
Following the similar protocol described above for the arithmetic question, we computed the entropy landscape, $S(L_{train},L_{test})$, as a function of train loss, $L_{train}$, and test loss, $L_{test}$. Results are presented in Figure 2 (a), where both $L_{train}$ and $L_{test}$ are on a log scale for clarity in the low-loss regime. In statistical physics, the probability distribution over different states (different locations in the plot) is sharply peaked around the entropy maximum, making the max-entropy state thermodynamic equilibrium. Therefore, we locate these max-entropy states at each $L_{train}$ (magenta dots), and compare with SGD trained states (black curve). SGD training results are averaged over 100 independent training instances using optimized hyperparameters, yielding error bars smaller than the black curve’s width. Our results suggest that at a given $L_{train}$, the corresponding max entropy loss is clearly lower than the test loss obtained via SGD training, demonstrating the high-entropy advantage in NNs generalizability. More technical details, including data and hyperparameter tuning, are presented in Supplementary Materials.
Figure 2: Entropy landscape of (a) Kaggle Housing Price dataset; (b) selected MNIST dataset; (c) Polymer SMILES dataset. For the SGD training of all three models, the standard error at each train loss level is smaller than the width of the curve.
@@
Modern neural networks have demonstrated remarkable success in computer vision.[11] One of the fundamental datasets used in the computer vision benchmark is the MNIST dataset, which consists of images with handwritten digits. [29] Here we used it to further verify that the high-entropy advantage for the computer vision task. Since many models have achieved 100% accuracy on the test set of MNIST (see Leaderboard of Kaggle Digit Recognizer competition[30]), we strategically increase the difficulty of the MNIST task by a 200 times reduction of the training dataset size. This is done by randomly sampling 500 images from the original dataset and dividing them equally as the train and test sets, respectively. We refer this dataset as the selected MNIST dataset in this paper. The increased difficulty allows us to better observe the potential advantage of the high-entropy states. We built a small convolutional neural network with 5 convolutional layers followed by one fully-connected layer. This network has 362 trainable parameters, which is larger than the training dataset (250 images) in the selected MNIST, making it overparamterized.
In Figure 2 (b), we presented the entropy landscape, $S(L_{train},A^{s}_{test})$, of this neural network on the selected MNIST dataset. Here $A^{s}_{test}$ is smoothed accuracy which is differentiable, as required by the WLMD algorithm (See Methods for details). We then find the max-entropy spot for each train loss on the landscape, which is again plotted as the magenta dots. Similar to the landscape in Figure 1 (a), we observed an increase of max entropy test accuracy when the train loss reaches a low level ($\ln(L_{train})\approx-1$) and saturation at $\ln(L_{train})\approx-4$. When the train loss is high, we observed a high-entropy band around the test accuracy of 0.1. This also meets our expectation because random guesses for this 10-class classification problem have an accuracy of 10%. We then performed classical training via the SGD optimizer for 200 times. (See more details in Supplementary Materials.) The mean SGD accuracy are plotted as the black curve. Our results suggest that for a given train loss, the corresponding max entropy accuracy generally outperforms the SGD training, especially when the train loss is small ($\ln(L_{train})<-2$).
In order to confirm the existence of the entropy advantage in deeper neural networks, we investigated a 10-layer ResNet [11] (43604 trainable parameters) on a selected CIFAR-10 dataset (5000 images, equally divided for training and testing). The results, presented in the Supplementary Materials, exhibit strong high-entropy advantage as well.
@@
Natural language modeling is another domain where neural networks have consistently outperformed traditional machine learning methods. [2, 3, 4] Language modeling is also inherently differs from other machine learning tasks due to its reliance on semantic complexity and contextuality. [31] Therefore, we further extended our tests to a language modeling task. Specifically, we focused on a new emerging class of models called Chemical Language Models, which are language models trained on large databases of SMILES (Simplified Molecular Input Line Entry System) [32] strings and present promising performance in both predictive and generative tasks.[33, 34] We utilized the TransPolymer model published recently, which achieves state-of-the-art performance in all ten different downstream tasks for polymer property prediction.[35] TransPolymer is a BERT (Bidirectional Encoder Representations from Transformers) family model[3, 36] and is pretrained on roughly 5 millions polymer SMILES augmented from the Pl1M database.[37] This pretrained transformer-based language model is able to take the SMILES as input directly and generate a 768-dimension embedding for each given SMILES. The embedding is then fed into a regressor head, a fully-connected layer with SiLU activation function, and is regressed on different polymer properties. Here we choose the Egb Dataset[38, 35] to perform the entropy advantage experiment, which is the bandgap energy of bulk polymer and consists of 561 data points. This is also the dataset where TransPolymer has the best performance (test $R^{2}=0.93$), therefore, we can verify whether entropy advantage still exists for such a well-learned task. 80% data is used for training and the remaining 20% data is reserved for testing, which is same as the way reported in the original TransPolymer paper.[35]
For sampling efficiency, we follow the fintuning strategy of large foundation model,[39, 40, 35] where the encoder embedding is fixed and only tuning the regressor head of the TransPolymer model. We also reduced the width of the regressor to 50 while keeping the SiLU activation function. Therefore, the final model has $38501$ parameters, making it overparamterized. Entropy landscape of this language modeling task is presented in Figure 2 (c). The SGD training curve is obtained by performing the classical training for 40 times. As a regression problem, we found that the corresponding max-entropy loss at each $L_{train}$ is slightly lower or comparable to its SGD training analog, which suggests that high-entropy state can generalize well. In other words, even for a task that could be well-learned via the SGD training, there is still a high-entropy advantage.
@@
Now that we have demonstrated high-entropy advantage in four distinct machine learning tasks, we will next study how this advantage is affected by the size of the neural network. It is known that neural networks are equivalent to Gaussian processes in the limit of infinite width.[41, 26] Therefore, it is reasonable to expect that the high-entropy advantage could also vary with the network width.
To better evaluate this, we constructed a Spiral Regression Dataset (see Supplementary Materials for details) which consists of 500 data points and is divided equally for training and test sets. Using the same WLMD method, we sampled the entropy landscapes on NNs with 2 hidden layers of four different widths, $W$, ranging from 30 to 1000. All four models are overparamerterized and the largest one has more than 1 million trainable parameters. Our results, presented in Fig. 3, suggest that the high-entropy advantage decreases as $W$ increases, and finally fades away when $W=1000$. Note here we used the Adam optimizer here because the SGD optimizer performs much worse. We also performed similar experiments on the Kaggle House price prediction, the selected MNIST image recognition, and the Polymer SMILES language modeling tasks. These additional experiments confirmed a similar trend: the high-entropy advantage is more significant for narrower networks (see Figure 7 - 9 in the Supplementary Materials).